A company recently bought a new robot with a first cost of $240,000, and a depreciation lifetime of five years. Using a table for the Modified Accelerated Cost Recovery System, please determine the depreciation amounts and book values for years 1-6. Also, if the robot is sold at the end of year 4 for $100,000, what would be the capital gain, depreciation recapture, or capital loss?
Answer #1

MACRS depreciation rates for six years are 0.20, 0.32, 0.1920, 0.1152, 0.1152 & 0.0576

Depreciation for year 1 = 0.20 * 240000 = 48000

BV at end of year 1 = 240000 - 48000 = 192000

Depreciation of year 2 = 240000 * 0.32 = 76800

BV at end of year 2 = 192000 - 76800 = 115200

Depreciation of year 3 = 240000 * 0.1920 = 46080

BV at end of year 3 = 115200 - 46080 = 69120

Depreciation of year 4 = 240000 * 0.1152 = 27648

BV at end of year 4 = 69120 - 27468 = 41472

Depreciation of year 5 = 240000 * 0.1152 = 27648

BV at end of year 5 = 41472 - 27468 = 13824

Depreciation of year 6 = 240000 * 0.0576 = 13824

BV at end of year 6 = 13824 - 13824 = 0

Capital gain = 100000 - 41472 = 58528